OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -7, an = 4 • an-1

OpenStudy (anonymous):

\[a _{n} = 4 • a _{n-1}\] Second Term \[a _{2} = 4 • a _{2-1}\] \[a _{2} = 4 • a _{1}\] \[a _{2} = 4 •(-7)\] \[a _{2} = -28\]

OpenStudy (anonymous):

\[a _{n} = 4 • a _{n-1}\] Third Term \[a _{3} = 4 • a _{3-1}\] \[a _{3} = 4 • a _{2}\] \[a _{3} = 4 •(-28)\] \[a _{3} = -112\]

OpenStudy (anonymous):

now find the other three terms :)

OpenStudy (anonymous):

@Ambition do you understand?

OpenStudy (anonymous):

yes -448, -1792, -7168

OpenStudy (anonymous):

Good Job! :D